# AGFS-Tractometry: A Novel Atlas-Guided Fine-Scale Tractometry Approach for Enhanced Along-Tract Group Statistical Comparison Using Diffusion MRI Tractography

### v1

@@ Abstract
00footnotetext: *Fan Zhang and Lauren J. O’Donnell are co-senior-authors.
Keywords: Diffusion MRI, Tractography, Tractometry, Group Comparison, Permutation Test.
@@ 1 Introduction
Diffusion magnetic resonance imaging (dMRI) is currently the only technique for in vivo reconstruction of the brain’s white matter (WM) connections, providing a noninvasive opportunity to characterize their underlying tissue microstructure by probing water molecule diffusion and mapping fiber tracts [1]. In tractography, fiber tracts are modeled as collections of streamlines that reflect the pathways of WM fibers traversing various anatomical regions. For example, Figure 1a gives a visualization of the corticospinal tract (CST) and its somatotopic organization, with four major subdivisions connecting to different cortical regions. Tractometry is an advanced tractography analysis technique for along-tract analysis to investigate the morphology and microstructural properties along anatomical fiber tracts [2, 3, 4, 5, 6]. Unlike traditional approaches that compute a summary statistic of the entire tract (e.g., streamline count or volume), tractometry enables the study of local regions along the fiber tract and thus can be more sensitive to detect subtle WM group differences for between-population comparisons [7]. Over the past decade, group comparison of the WM fiber tracts across populations using tractometry has increasingly contributed to our understanding of normal brain development and assisted the identification of potential biomarkers for neurological disorders [2, 8, 9].
@@
In general, the first step of performing tractometric analyses is tract profiling, which describes the distribution of the WM measures along the fiber tract. In this step, the fiber tract needs to be segmented into smaller parcels, from which the WM measure of interest is extracted (e.g., average fractional anisotropy, FA). Currently, state-of-the-art methods for tract profiling are the AFQ [4] and BUAN [5] methods. In AFQ, streamlines are resampled to $n$ equidistant points along their arc length (e.g., $n=100$), and each tract parcel is determined by its position indices of points along the streamline (Figure 1b, left). BUAN proposes a more advanced parcellation approach to assign each streamline point to the closest point on a tract-centerline composed of n points (Figure 1b, right). Despite the widespread use of these two methods, a limitation is their focus on the entire fiber tract, overlooking possible anatomical subdivisions within the tract. For example, the CST has a somatotopic organization with multiple subdivisions connecting to the leg, trunk, hand, and face cortical areas [10] (see Figure 1a). The existing AFQ and BUAN methods are not effective in distinguishing tract parcels belonging to the different anatomical subdivisions. Therefore, we propose a novel tract profiling strategy that leverages a fine-scale along-tract parcellation to enable a detailed investigation of local tract regions (Figure 1c).
@@
After creating the tract profile, the next step in tractometry is statistical comparison, which aims to identify local tract regions with statistically significant group differences. Current tractometry methods typically apply hypothesis testing at each parcel along the tract profile. For example, AFQ uses Student’s t-test at each node of the tract profile, whereas BUAN employs a linear mixed model (LMM) to account for both fixed and random effects in the experimental data. In both cases, multiple comparisons across multiple tract parcels are controlled using methods such as the false discovery rate (FDR) and permutation test [11]. The large number of tract parcels in tractometric analyses, especially in fine-scale profiling, increases the number of statistical tests, which in turn demands stricter correction for multiple comparisons. As a result, the aforementioned conventional statistical analysis methods may lack the sensitivity required to reliably detect group differences. This limitation is a well-recognized challenge in voxel-based neuroimaging approaches such as tract-based spatial statistics (TBSS) [12] and voxel-based morphometry (VBM) [13], which involve conducting a large number of statistical tests across individual voxels and require correction for multiple comparisons. Currently, clustering-thresholding techniques are widely used in neuroimaging to enhance statistical sensitivity by leveraging spatial contiguity, particularly in voxel-based analysis (VBA) [14, 15]. In tractography, similar approaches have been applied to detect tract-level WM differences [16, 17] or network changes [18], whereas no prior work has applied cluster-thresholding in tractometry. In our study, we propose to use clustering-thresholding-based statistics for tractometry group analysis, where multiple along-tract comparisons present unique statistical challenges.
@@
In this paper, we design a novel atlas-guided fine-scale tractometry method, namely AGFS-Tractometry, that leverages tract spatial information and permutation testing to enhance the along-tract statistical analysis between populations. Our study has two major contributions, corresponding to the two major steps in tractometric analyses. First, for tract profiling, we introduce a novel approach that begins with the creation of a population-wise tract profiling template. This template is constructed based on an anatomically curated WM fiber tract atlas (i.e., the ORG-atlas) [19, 20], where each anatomical fiber tract is divided into multiple fiber clusters. This allows us to derive along-tract parcellations from each cluster, resulting in a finer-scale parcellation of the tract of interest (as illustrated in Figure 1). Furthermore, the created template can be applied to individual tractography data to extract subject-specific tract profiles. In this way, using the template acts as a common space where all subjects are mapped into a unified framework. This naturally establishes correspondence for local along-tract parcels across multiple subjects, facilitating robust group comparisons. Second, for group-wise statistical comparison, we introduce a novel nonparametric, permutation testing method to enable simultaneous analysis across all along-tract parcels while correcting for multiple comparisons. The design of this method is inspired by voxel-based cluster-thresholding approaches, which leverage spatial neighborhoods to strengthen statistical sensitivity [14]. In our approach, we develop a parcel spatial neighborhood construction strategy to form spatially adjacent parcels into communities. Then the parcels within communities can mutually reinforce each other to enhance the detection of statistically significant differences. We evaluate our AGFS-Tractometry by comparing it with two widely used tractometry techniques, AFQ and BUAN. First, we conduct experiments on synthetic datasets that include ground truth, meaning the group differences along the fiber tracts are known. Our approach exhibits enhanced sensitivity and specificity in detecting these ground truth differences. Additionally, we assess our method in two real-world scenarios: one focuses on investigating sex differences in WM fiber tracts, and the other on examining the effects of disease on WM in autism. The results demonstrate that our method achieves improved sensitivity in identifying localized differences.
@@ 2.2.1 Fine-Scale Along-Tract Parcellation
Specifically, in the ORG atlas, a fiber cluster is represented as a collection of streamlines, where each streamline consists of multiple points along its length. For each cluster, we adopt the centerline profiling strategy proposed in BUAN for along-cluster parcellation. A centerline is first generated by resampling each fiber streamline to n equidistant points and computing the average coordinates of the corresponding points across all streamlines at each location. Next, the spatial distances between each streamline point in the cluster and each centerline point are calculated, followed by assigning each streamline point to its nearest centerline point. This process results in an along-cluster parcellation of each fiber cluster into n parcels. One advantage of performing cluster-level centerline profiling is that in the ORG atlas, fiber clusters are created by grouping streamlines that share similar geometric trajectories and lengths. In this way, streamline points within the same parcel maintain comparable distances to the centerline point, ensuring that each parcel reflects the local geometry of fiber clusters. This can avoid a potential issue observed in BUAN, where imperfect alignment at tract endpoint regions may lead to the misclassification of points from the same location into different parcels (see Figure 1b). Lastly, for a specific anatomical tract composed of $m$ clusters, a fine-scale along-tract parcellation into $m$ $n$ parcels is obtained to serve as the template for tract profiling.
@@ 2.3.2 Quantitative Measure Computation
where $D_{M}(\mathbf{x},\boldsymbol{\mu})$ represents the Mahalanobis distance between a streamline point $\mathbf{x}$ and the mean $\boldsymbol{\mu}$ of a distribution with covariance matrix $\boldsymbol{\Sigma}$ computed by all points within the parcel. $\mathrm{Metric}_{i}$ is the diffusion metric for the $i$-th point (e.g., FA). The parcel-level diffusion metric, $\mathrm{Metric}_{\mathrm{weighted}}$, is then computed as a weighted average, where the weights are derived from the Mahalanobis distance $D_{M}(\mathbf{x},\boldsymbol{\mu})$. In our study, the diffusion measure FA is used, while we note that any other measures of interest (e.g., mean diffusivity (MD) and number of fiber points) can be used for quantitative parcel measures to compute the subject-specific tract profile.
@@ 2.4.1 Parcel-Level Statistical Comparison
A parcel-level statistical test is performed for each along-tract parcel to identify the parcels with uncorrected group differences at a predetermined primary threshold. These parcels are referred to as the suprathreshold parcels. Specifically, we perform a null hypothesis test that computes group differences for each parcel to derive an uncorrected p-value. Commonly used statistical test methods, such as Student’s t-test and generalized linear model (GLM), can be used. We set the primary threshold to p-value$=0.05$, i.e., the parcels with uncorrected p-values $\leq 0.05$ pass the test.
@@ 2.4.2 Suprathreshold Parcel Community Formation
Using the parcel neighborhood graph derived from all along-tract parcels (as computed using the ORG atlas prior to statistical analysis; see Section 2.2.1), we first construct two $n$×$n$ binary matrices: (1) a parcel neighborhood matrix indicating whether each pair of parcels are spatially adjacent (Section 2.2.2), and (2) a parcel-level significance matrix, where an element is 1 only if a pair of parcels are both suprathreshold in the statistical testing (Section 2.4.1). We then perform an element-wise matrix multiplication between the neighborhood matrix and the parcel-level significance matrix. This operation retains only the neighborhood graph edges connecting suprathreshold parcels, effectively removing edges between non-significant nodes, and thus retaining subgraphs of connected suprathreshold parcels. Based on the retained subgraphs, CP is applied to identify all cliques and then integrate the adjacent parcel cliques as communities, as illustrated in Figure 4. In our experiments, we set $k$=3, $h$=2 that defined a pair of 3-cliques ($k$ = 3) as adjacent if there were at least two shared parcels ($h$ = 2). We note that the parameter selections of $k$=3 and $h$=2 are the minimum values that allow for CP [26], ensuring all possible suprathreshold parcel communities are identified.
@@ 3.1.2 Whole-Brain Tractography
Whole-brain tractography is performed using the UKF tractography method [23, 24, 25], as implemented in the ukftractography package via SlicerDMRI [36, 37]. The UKF approach fits a two-tensor mixture model to the dMRI data while tracking fibers, incorporating prior information from the previous step to enhance model stability. Previous studies have demonstrated that UKF accounts for crossing fibers and offers sensitive and reliable fiber tracking with a wide range of populations, including variations in age, health conditions, and dMRI acquisition protocols [19, 38, 39]. Furthermore, UKF performs simultaneous fiber tracking and multi-tensor modeling so that the microstructure measures (e.g., FA and MD) can be directly computed at points along each fiber streamline. This allows the estimation of fiber-specific microstructure estimation for accurate parcel measure computation. In our study, for each of the dMRI scans from the individual subjects, tractography parameters are adopted from previous work [19] for whole-brain WM fiber tracking. Specifically, tractography was seeded in all voxels within the brain mask where FA exceeded 0.1. Tracking was terminated when FA dropped below 0.08 or the normalized mean signal decreased below 0.06. The diffusion tensors, along with their FA values, were preserved for subsequent analyses, facilitating a comprehensive assessment of the WM microstructure.
@@ 3.4 Real Data Analysis
In the interest of the application to study inter-group significant differences in brain WM across populations, we apply our proposed method to two different datasets: 1) the HCP-YA dataset and 2) the ABIDE II-NYU dataset. The HCP dataset includes 100 healthy adult subjects, consisting of 54 females and 46 males. We conduct a study to investigate sex differences in the brain WM using this dataset. The ABIDE dataset includes 48 individuals with ASD and 22 HC individuals. This dataset is used to study ASD-related WM differences compared to HC.
@@ 3.4.1 Sex-related WM Difference Using the HCP-YA Dataset
There is an evident prevalence of sex-specific differences in the brain WM, which has greatly advanced our understanding of behavioral variations between females and males [40, 41]. In particular, language-related aspects are regarded as a critical factor in explaining sex-based behavioral differences [42, 43, 44], with extensive research demonstrating their strong association with brain WM [45, 46]. The AF is primarily associated with language functions, connecting the frontal and temporal lobes, and plays a crucial role in language processing mechanisms, such as speech production, auditory comprehension, and verbal memory [47, 48]. It links Broca’s area to Wernicke’s area, facilitating communication between regions involved in language comprehension and production [49]. In the interest of our study, we apply the three compared tractometry techniques (AFQ, BUAN, and our AGFS-Tractometry) to assess their ability in detecting sex-specific differences in local regions of the AF tract.

### v2

@@ Abstract
keywords:
Diffusion MRI , Tractography , Tractometry , Group Comparison , Permutation Test.
journal: Medical Image Analysis
[…]
[a]organization=University of Electronic Science and Technology of China, addressline=No. 2006, Xiyuan Ave, city=Chengdu, postcode=611731, state=Sichuan, country=China
[…]
[b]organization=Brigham and Women’s Hospital, Harvard Medical School, addressline=75 Francis St, city=Boston, postcode=02215, state=MA, country=USA
[…]
[c]organization=Boston University, addressline=72 E Concord St, city=Boston, postcode=02118, state=MA, country=USA
@@ 1 Introduction
Diffusion magnetic resonance imaging (dMRI) is currently the only technique for in vivo reconstruction of the brain’s white matter (WM) connections, providing a noninvasive opportunity to characterize their underlying tissue microstructure by probing water molecule diffusion and mapping fiber tracts [2]. In tractography, fiber tracts are modeled as collections of streamlines that reflect the pathways of WM fibers traversing various anatomical regions. For example, Figure 1a gives a visualization of the corticospinal tract(CST [^1] In line with common practice in the tractography field, the term ”CST” is used throughout to describe the major descending pathway from the motor cortex. We acknowledge that the CST is a subpart of the PyT, which also includes several other fibers projecting to deep GM structures[30]., i.e., its trajectory within the pyramidal tract/PyT) [30] and its somatotopic organization, with four major subdivisions connecting to different cortical regions. Tractometry is an advanced tractography analysis technique for along-tract analysis to investigate the morphology and microstructural properties along anatomical fiber tracts [11, 12, 60, 9, 41]. Unlike traditional approaches that compute a summary statistic of the entire tract (e.g., streamline count or volume), tractometry enables the study of local regions along the fiber tract and thus can be more sensitive to detect subtle WM group differences for between-population comparisons [66]. Over the past decade, group comparison of the WM fiber tracts across populations using tractometry has increasingly contributed to our understanding of normal brain development and assisted the identification of potential biomarkers for neurological disorders [11, 8, 36].
@@
In general, the first step of performing tractometric analyses is tract profiling, which describes the distribution of the WM measures along the fiber tract. In this step, the fiber tract needs to be segmented into smaller parcels, from which the WM measure of interest is extracted (e.g., average fractional anisotropy, FA). Currently, state-of-the-art methods for tract profiling are the AFQ [60] and BUAN [9] methods. In AFQ, streamlines are resampled to $n$ equidistant points along their arc length (e.g., $n=100$), and each tract parcel is determined by its position indices of points along the streamline (Figure 1b, left). BUAN proposes a more advanced parcellation approach to assign each streamline point to the closest point on a tract-centerline composed of n points (Figure 1b, right). Despite the widespread use of these two methods, a limitation is their focus on the entire fiber tract, overlooking possible anatomical subdivisions within the tract. For example, the CST pathway has a somatotopic organization with multiple subdivisions connecting to the leg, trunk, hand, and face cortical areas [22] (see Figure 1a). The existing AFQ and BUAN methods are not effective in distinguishing tract parcels belonging to the different anatomical subdivisions. Therefore, we propose a novel tract profiling strategy that leverages what we term a fine-scale along-tract parcellation. This fine-scale approach provides a higher spatial sampling resolution by parcellating the tract not only along its length, as in AFQ and BUAN, but also across its diameter by leveraging fiber cluster subdivisions. This enables a detailed investigation of local tract regions (Figure 1c) and and yields a substantially greater number of parcels than conventional approaches (e.g., increasing from 100 parcels along the tract in AFQ to 100 parcels per cluster × 15 clusters = 1500 parcels in total).
@@
After creating the tract profile, the next step in tractometry is statistical comparison, which aims to identify local tract regions with statistically significant group differences. Current tractometry methods typically apply hypothesis testing at each parcel along the tract profile. For example, AFQ uses Student’s t-test at each node of the tract profile, whereas BUAN employs a linear mixed model (LMM) to account for both fixed and random effects in the experimental data. In both cases, multiple comparisons across multiple tract parcels are controlled using methods such as the false discovery rate (FDR) and permutation test [3]. The large number of tract parcels in tractometric analyses, especially in fine-scale profiling, increases the number of statistical tests, which in turn demands stricter correction for multiple comparisons. As a result, the aforementioned conventional statistical analysis methods may lack the sensitivity required to reliably detect group differences. This limitation is a well-recognized challenge in voxel-based neuroimaging approaches such as tract-based spatial statistics (TBSS) [49] and voxel-based morphometry (VBM) [1], which involve conducting a large number of statistical tests across individual voxels and require correction for multiple comparisons. Currently, clustering-thresholding techniques are widely used in neuroimaging to enhance statistical sensitivity by leveraging spatial contiguity, particularly in voxel-based analysis (VBA) [37, 56]. In tractography, similar approaches have been applied to detect tract-level WM differences [68, 57] or structural network changes [64]. In the context of tractometry, cluster-wise correction has been explored in approaches such as correlational tractography [61] and Tracula [62], which leverage spatial clustering strategies to improve sensitivity. However, no prior work has applied cluster-thresholding in tractometry, where multiple along-tract comparisons pose a statistical challenge. In our study, we introduce a novel tractometry framework for anatomically guided fiber profiling and propose a clustering-thresholding-based statistical method to enable simultaneous analysis across all along-tract parcels while correcting for multiple comparisons.
@@
In this paper, we design a novel atlas-guided fine-scale tractometry method, namely AGFS-Tractometry, that leverages tract spatial information and permutation testing to enhance the along-tract statistical analysis between populations. Our study has two major contributions, corresponding to the two major steps in tractometric analyses. First, for tract profiling, we introduce a novel approach that begins with the creation of a population-wise tract profiling template. This template is constructed based on an anatomically curated WM fiber tract atlas (i.e., the ORG-atlas) [70, 69]. In this atlas, major tracts are subdivided into multiple fiber clusters. Our method leverages these clusters to achieve a fine-scale parcellation, which we define as a higher spatial sampling resolution achieved by parcellating not only along the tract’s length but also across its diameter (as illustrated in Figure 1). Furthermore, the created template can be applied to individual tractography data to extract subject-specific tract profiles. In this way, using the template acts as a common space where all subjects are mapped into a unified framework. This naturally establishes correspondence for local along-tract parcels across multiple subjects, facilitating robust group comparisons. Second, for group-wise statistical comparison, we introduce a novel nonparametric, permutation testing method to enable simultaneous analysis across all along-tract parcels while correcting for multiple comparisons. The design of this method is inspired by voxel-based cluster-thresholding approaches, which leverage spatial neighborhoods to strengthen statistical sensitivity [37]. In our approach, we develop a parcel spatial neighborhood construction strategy to form spatially adjacent parcels into communities. Then the parcels within communities can mutually reinforce each other to enhance the detection of statistically significant differences. We evaluate our AGFS-Tractometry by comparing it with two widely used tractometry techniques, AFQ and BUAN. First, we conduct experiments on synthetic datasets that include ground truth, meaning the group differences along the fiber tracts are known. Our approach exhibits enhanced sensitivity and specificity in detecting these ground truth differences. Additionally, we assess our method in two real-world scenarios: one focuses on investigating sex differences in WM fiber tracts, and the other on examining the effects of disease on WM in autism. The results demonstrate that our method achieves improved sensitivity in identifying localized differences.
@@ 2.2.1 Fine-Scale Along-Tract Parcellation
Specifically, in the ORG atlas, a fiber cluster is represented as a collection of streamlines, where each streamline consists of multiple points along its length. For each cluster, we adopt the centerline profiling strategy proposed in BUAN for along-cluster parcellation. A centerline is first generated by resampling each fiber streamline to n equidistant points and computing the average coordinates of the corresponding points across all streamlines at each location. Next, the spatial distances between each streamline point in the cluster and each centerline point are calculated, followed by assigning each streamline point to its nearest centerline point. This process results in an along-cluster parcellation of each fiber cluster into n parcels. One advantage of performing cluster-level centerline profiling is that in the ORG atlas, fiber clusters are created by grouping streamlines that share similar geometric trajectories and lengths. In this way, streamline points within the same parcel maintain comparable distances to the centerline point, ensuring that each parcel reflects the local geometry of fiber clusters. This can avoid a potential issue observed in BUAN, where imperfect alignment at tract endpoint regions may lead to the misclassification of points from the same location into different parcels (see Figure 1b). Lastly, for a specific anatomical tract composed of $m$ clusters, a fine-scale along-tract parcellation into $m$ × $n$ parcels is obtained to serve as the template for tract profiling.
@@ 2.3.2 Quantitative Measure Computation
where $D_{M}(\mathbf{x},\boldsymbol{\mu})$ represents the Mahalanobis distance between a streamline point $\mathbf{x}$ and the mean $\boldsymbol{\mu}$ of a distribution with covariance matrix $\boldsymbol{\Sigma}$ computed by all points within the parcel. $\mathrm{Metric}_{i}$ is the diffusion metric for the $i$-th point (e.g., FA). The parcel-level diffusion metric, $\mathrm{Metric}_{\mathrm{weighted}}$, is then computed as a weighted average, where the weights are derived from the Mahalanobis distance $D_{M}(\mathbf{x},\boldsymbol{\mu})$. In our study, the diffusion measure FA is used, while we note that any other measures of interest (e.g., mean diffusivity (MD) and number of fiber points) can be used for quantitative parcel measures to compute the subject-specific tract profile. Furthermore, it is important to note that while the UKF is employed in this study for fiber tracking, our proposed framework is not limited to this specific method. The framework is designed to be generalizable and can accommodate any fiber tracking algorithm that provides point-specific microstructural measures along fiber streamlines. This flexibility allows for the integration of various fiber tracking techniques, ensuring that our method can be adapted to different research contexts and data types.
@@ 2.4.1 Parcel-Level Statistical Comparison
A parcel-level statistical test is performed for each along-tract parcel to identify the parcels with uncorrected group differences at a predetermined primary threshold. These parcels are referred to as the suprathreshold parcels. Specifically, we perform a null hypothesis test that computes group differences for each parcel to derive an uncorrected p-value. Commonly used statistical test methods, such as Student’s t-test and generalized linear model (GLM), can be used. We set the primary threshold to p-value=0.05=0.05, i.e., the parcels with uncorrected p-values $\leq 0.05$ pass the test.
@@ 2.4.2 Suprathreshold Parcel Community Formation
Using the parcel neighborhood graph derived from all along-tract parcels (as computed using the ORG atlas prior to statistical analysis; see Section 2.2.1), we first construct two nn×$n$ binary matrices: (1) a parcel neighborhood matrix indicating whether each pair of parcels are spatially adjacent (Section 2.2.2), and (2) a parcel-level significance matrix, where an element is 1 only if a pair of parcels are both suprathreshold in the statistical testing (Section 2.4.1). We then perform an element-wise matrix multiplication between the neighborhood matrix and the parcel-level significance matrix. This operation retains only the neighborhood graph edges connecting suprathreshold parcels, effectively removing edges between non-significant nodes, and thus retaining subgraphs of connected suprathreshold parcels. Based on the retained subgraphs, CP is applied to identify all cliques and then integrate the adjacent parcel cliques as communities, as illustrated in Figure 4. In our experiments, we set kk=3, hh=2 that defined a pair of 3-cliques ($k$ = 3) as adjacent if there were at least two shared parcels ($h$ = 2). We note that the parameter selections of kk=3 and hh=2 are the minimum values that allow for CP [43], ensuring all possible suprathreshold parcel communities are identified.
@@ 3.1.2 Whole-Brain Tractography
Whole-brain tractography is performed using the UKF tractography method [45, 54, 16], as implemented in the ukftractography package via SlicerDMRI [38, 39]. The UKF approach fits a two-tensor mixture model to the dMRI data while tracking fibers, incorporating prior information from the previous step to enhance model stability. Previous studies have demonstrated that UKF accounts for crossing fibers and offers sensitive and reliable fiber tracking with a wide range of populations, including variations in age, health conditions, and dMRI acquisition protocols [70, 65, 59]. Furthermore, UKF performs simultaneous fiber tracking and multi-tensor modeling. This is more robust to partial volume effects (PVEs), allowing microstructure measures (including FA) to be directly computed at points along each fiber streamline, enabling fiber-specific microstructure estimation for accurate parcel measure computation. However, it is critical to distinguish between the pathway identified by tractography and the biological interpretation of its microstructure. For example, while we can delineate the trajectory of the CST pathway, the microstructural metrics (e.g., FA) computed along this pathway inherently reflect the average properties of the entire PyT. This is because the spatial resolution of dMRI cannot separate CST axons from other intermingled corticofugal fibers (e.g., corticobulbar fibers) within the PyT at the sub-voxel level [30].
In our study, for each of the dMRI scans from the individual subjects, tractography parameters are adopted from previous work [70] for whole-brain WM fiber tracking. Specifically, tractography was seeded in all voxels within the brain mask where FA exceeded 0.1. Tracking was terminated when FA dropped below 0.08 or the normalized mean signal decreased below 0.06. The diffusion tensors, along with their FA values, were preserved for subsequent analyses, facilitating a comprehensive assessment of the WM microstructure.
@@ 3.4 Real Data Analysis
In the interest of the application to study inter-group significant differences in brain WM across populations, we apply our proposed method to two different datasets: 1) the HCP-YA dataset and 2) the ABIDE II-NYU dataset. The HCP dataset includes 100 healthy adult subjects, consisting of 54 females and 46 males. We conduct a study to investigate sex differences in the brain WM using this dataset. Additionally, to assess the reproducibility of the identified sex differences, we performed a test-retest replication analysis using another 100 subjects from the same HCP-YA dataset. This replication cohort maintained the same sex composition (54 females and 46 males) and showed no significant difference in age distribution. The ABIDE dataset includes 48 individuals with ASD and 22 HC individuals. This dataset is used to study ASD-related WM differences compared to HC.
@@ 3.4.1 Sex-related WM Difference Using the HCP-YA Dataset
There is an evident prevalence of sex-specific differences in the brain WM, which has greatly advanced our understanding of behavioral variations between females and males [13, 19]. In particular, language-related aspects are regarded as a critical factor in explaining sex-based behavioral differences [48, 58, 21], with extensive research demonstrating their strong association with brain WM [44, 33]. The AF is primarily associated with language functions, connecting the frontal and temporal lobes, and plays a crucial role in language processing mechanisms, such as speech production, auditory comprehension, and verbal memory [5, 7]. It links Broca’s area to Wernicke’s area, facilitating communication between regions involved in language comprehension and production [34]. In the interest of our study, we apply the three compared tractometry techniques (AFQ, BUAN, and our AGFS-Tractometry) to assess their ability in detecting sex-specific differences in local regions of the AF tract. For AFQ, we adhered to the standard practice of using Student’s t-test, as preliminary analyses indicated that applying a LMM was not feasible due to insufficient data points per subject for reliable random effects estimation.
@@
To further validate the robustness and reproducibility of our findings, we conducted two additional experiments: (1) a test-retest reliability analysis and (2) an independent cohort validation. The details are as follows:
Test-Retest Reliability: We evaluated the test-retest reliability of our proposed AGFS-Tractometry method using the HCP test-retest dataset (N=36 subjects). For the 11 clusters identified along the left AF, we computed the intraclass correlation coefficient (ICC) between the test and retest scans for each corresponding parcel. The results demonstrated good to excellent reliability across all clusters, with 4 clusters exhibiting ICC values above 0.75 and the remaining 7 clusters falling in the range of 0.60-0.74. According to the established guidelines by (Cicchetti 1994), this indicates a high degree of measurement consistency. Furthermore, a direct group comparison between the test and retest data using our method revealed no statistically significant differences in the left AF, confirming the high reproducibility of our measurements.
Independent Cohort Validation: To demonstrate the generalizability of our initial sex-difference findings, we replicated the analysis on an independent, matched cohort of 100 new subjects from the HCP dataset. This new cohort was selected to have an identical sex distribution and a comparable age range (with no significant difference in age) to the 100 original subjects. Applying our method to this validation cohort, we successfully identified localized sex-related differences in the left AF. The spatial pattern of these significant regions was highly consistent with our original results in Figure 6. Specifically, when the significant regions from both analyses were binarized and compared as volumetric masks, they exhibited an 84.24% spatial overlap, predominantly located near Broca’s and Wernicke’s areas. This strong concordance across independent datasets underscores the robustness and reliability of our proposed tractometry framework.